• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Holmstrom

    Log entry time 10:26:04 on July18,2003

    Entry number 104468

    keyword=Set Septum Current to 310A

    We have set the septum current to 310 A, so that target tests can be done.